Will a precipitate form if the pH of a 0.015 M solution of Ni(NO3)2 is adjusted to 9.00? Ksp Ni(OH)2 = 1.7 x10-6 .
Answer #1

pH = 9.00

pOH = 14 - pH

pOH = 14 - 9.00

pOH = 5.00

[OH-] = 10-pOH

[OH-] = 10-5.00

[OH-] = 1.0 x 10-5 M

initial concentration Ni2+ = [Ni(NO3)2] = 0.015 M

Reaction quotient, Q = [Ni2+][OH-]2

Q = (0.015 M) * (1.0 x 10-5 M)2

Q = 1.5 x 10-12

Since Q < Ksp, therefore, precipitate will not form.